# Quantifying Classifier Utility under Local Differential Privacy

### v2

@@ C.1.2 Non-closed-form Classifiers
C.2 LDP Mechanisms in Figure 4
@@
C.2.1 Piecewise-based Mechanism
The original PM mechanism [54] is defined on $1,1]\to[-C_{\varepsilon},C_{\varepsilon}]$, where $C_{\varepsilon}$ is a variable dependent on $\varepsilon$. Such design ensures unbiasedness, as it was originally designed for mean estimation. The original SW mechanism [37] is defined on $0,1]\to[-b_{\varepsilon},1+b_{\varepsilon}]$. Unlike PM, SW is biased and designed for distribution estimation.
@@ C.2.1 Piecewise-based Mechanism
$\rho(\varepsilon,\theta)$curves. The concentration analysis $F(x+\theta)-F(x-\theta)$ of the PM and SW mechanisms results in a linear curve with two segments. When $\theta$ is small, the concentration is dominated by the $l_{x,\varepsilon},r_{x,\varepsilon})$ segment, which gives a slope proportional to $p_{\varepsilon}$. When $\theta$ is large, the concentration is dominated by the $0,1)\,\backslash\,[l_{x,\varepsilon},r_{x,\varepsilon})$ segment, which gives a slope proportional to $p_{\varepsilon}/\exp{(\varepsilon)}$. Thus, the $\rho(\varepsilon,\theta)$ curves for the PM and SW mechanisms exhibit a linear pattern with two distinct slopes.
C.2.2 Discrete Mechanism
@@
$\rho(\varepsilon,\theta)$curves. The concentration analysis $F(x+\theta)-F(x-\theta)$ of the Exponential mechanism results in a smooth curve, as the probability of seeing value $\tilde{x}$ further away from $x$ decreases smoothly w.r.t $\theta$. In contrast, the concentration analysis of the $k$-$R$ mechanism exhibits a pattern similar to the PM and SW mechanisms. When $\theta$ is small, the probability is dominated by the bin containing $x$, results in a slope $\propto\exp(\varepsilon)/(|\mathcal{X}-1+\exp(\varepsilon)|)$. When $\theta$ becomes larger, the probability is dominated by the other bins, which gives a slope $\propto 1/(|\mathcal{X}-1+\exp(\varepsilon)|)$.
C.3 PAC LDP vs $(\varepsilon,\delta)$-LDP (Section 5.2)
@@
Definition 15 ($(\varepsilon,\delta)$-LDP [17]).
@@
Given concrete values of pair $(\varepsilon,\delta)$, [^11] In fact, $\delta$ can be defined in a more intricate manner, such as incorporating the noise scale $\sigma$ as in [7]. However, this approach introduces a recursive dependency thus uncontrollable in practical applications, as $\sigma$ itself should be defined in terms of $\delta$. the term $\delta/\exp(\varepsilon)$ is a constant. Thus, it means that the distance between the distributions of $\mathcal{M}(x_{1})$ and $\mathcal{M}(x_{2})$ must be strictly bounded, i.e. similar to pure $\varepsilon$-LDP. However, this definition is often interpreted as: it is $\varepsilon$-LDP with a failure probability of $\delta$. Meanwhile, the Gaussian mechanism relies on this interpretation.
@@
Other relaxed privacy notions, including the Concentrated privacy from Dwork [15], Renyi privacy [46], and PAC privacy [62], provide patches to $(\varepsilon,\delta)$-LDP. The Gaussian mechanism has natural results under the Concentrated privacy and Renyi privacy notions, while we provide a PAC privacy result in Appendix B.2. Among these privacy notions, $(\varepsilon,\delta)$-PAC LDP provides the same meaning as the common interpretation of $(\varepsilon,\delta)$-LDP, i.e. it is $\varepsilon$-LDP with a failure probability of $\delta$.
C.4 More Case Studies for the Stroke Prediction Dataset (Section 7.2.1)
@@
Figure 9: Empirical and theoretical utility for two classifiers trained on the Stroke Prediction dataset.
C.4.1 Other Records and Mechanisms
@@
C.4.2 Closed Form CDF of the PM Mechanism
@@
C.4.3 Projected Decision Boundary
(a) Logistic Regression (LR).
(b) Random Forest (RF).
Figure 10: Projected decision boundary of the two classifiers on the stroke dataset.
@@ C.4.3 Projected Decision Boundary
C.4.4 Average-case and Worst-case Utility
@@
Conclusion on Robustness. From the results of average-case and worst-case utility analysis of the Logistic Regression and Random Forest classifiers, we can conclude that the Logistic Regression classifier is more robust under LDP perturbations than the Random Forest classifier.
(a) LR: Average-case utility.
(b) LR: Worst-case utility.
Figure 11: Average-case and worst-case utility for the Logistic Regression classifier trained on the Stroke Prediction dataset.
(a) RF: Average-case utility.
(b) RF: Worst-case utility.
Figure 12: Average-case and worst-case utility for the Random Forest classifier trained on the Stroke Prediction dataset.
C.5 More Case Studies for the Bank Customer Attrition Dataset (Section 7.2.2)
C.5.1 Detailed Quantification
@@ C.5.2 Average-case and Worst-case Utility
Figure 13: Average-case and worst-case utility for the Logistic Regression classifier trained on the Bank Customer Attrition dataset.
@@
Figure 14: Average-case and worst-case utility for the Random Forest classifier trained on the Bank Customer Attrition dataset.
C.6 More Case Studies for the MNIST-$\times7$ Dataset (Section 7.2.3)
C.6.1 Detailed Discussion
In practice, not all dimensions are sensitive and require privacy protection. While we treat all dimensions as sensitive in the MNIST-$\times7$ dataset, users can opt to apply LDP mechanisms to only a subset of dimensions, i.e. specific parts of the image. In such cases, the theoretical utility quantification is more accurate.
C.6.2 Average-case and Worst-case Utility
We use the first 50 images from the MNIST-$\times7$ dataset as representative samples for analyzing the average-case and worst-case utility under LDP mechanisms. Figure 16 presents the results of this analysis. Both the average-case and worst-case utility show a similar trend to the point-wise utility quantification, with the PM mechanism consistently achieving the highest utility.
(a) NN: Utility on the first image in the dataset.
(b) NN: Utility on the second image in the dataset.
Figure 15: (Same as Figure 8) Empirical and theoretical utility for a Neural Network classifier trained on the MNIST-$\times7$ dataset.
@@ C.6.2 Average-case and Worst-case Utility
Figure 16: Average-case and worst-case utility for the Neural Network classifier trained on the MNIST-$\times7$ dataset.

### v3

@@ C.1.2. Non-closed-form Classifiers
C.2. Explanation of Definition 5
[…]
C.3. LDP Mechanisms in Figure 4
@@
C.3.1. Piecewise-based Mechanism
The original PM mechanism (DBLP:conf/icde/WangXYZHSS019) is defined on $1,1]\to[-C_{\varepsilon},C_{\varepsilon}]$, where $C_{\varepsilon}$ is a variable dependent on $\varepsilon$. Such design ensures unbiasedness, as it was originally designed for mean estimation. The original SW mechanism (DBLP:conf/sigmod/Li0LLS20) is defined on $0,1]\to[-b_{\varepsilon},1+b_{\varepsilon}]$. Unlike PM, SW is biased and designed for distribution estimation.
@@ C.3.1. Piecewise-based Mechanism
$\rho(\varepsilon,\theta)$curves. The concentration analysis $F(x+\theta)-F(x-\theta)$ of the PM and SW mechanisms results in a linear curve with two segments. When $\theta$ is small, the concentration is determined by the $l_{x,\varepsilon},r_{x,\varepsilon})$ segment, which gives a slope proportional to $p_{\varepsilon}$. When $\theta$ is large, the concentration is also determined by the $0,1)\,\backslash\,[l_{x,\varepsilon},r_{x,\varepsilon})$ segment, which gives a slope proportional to $p_{\varepsilon}/\exp{(\varepsilon)}$. Thus, the $\rho(\varepsilon,\theta)$ curves for the PM and SW mechanisms exhibit a linear pattern with two distinct slopes.
C.3.2. Discrete Mechanism
@@
$\rho(\varepsilon,\theta)$curves. The concentration analysis $F(x+\theta)-F(x-\theta)$ of the Exponential mechanism results in a smooth curve, as the probability of seeing value $\tilde{x}$ further away from $x$ decreases smoothly w.r.t $\theta$. In contrast, the concentration analysis of the $k$-$R$ mechanism exhibits a pattern similar to the PM and SW mechanisms. When $\theta$ is small, the probability is determined by the bin containing $x$, results in a slope $\propto\exp(\varepsilon)/(|\mathcal{X}-1+\exp(\varepsilon)|)$. When $\theta$ becomes larger, the probability is also determined by the other bins, which gives a slope $\propto 1/(|\mathcal{X}-1+\exp(\varepsilon)|)$.
C.4. PAC LDP vs $(\varepsilon,\delta)$-LDP (Section 5.2)
@@
Definition 0 ($(\varepsilon,\delta)$-LDP (DBLP:journals/fttcs/DworkR14)).
@@
Given concrete values of pair $(\varepsilon,\delta)$, [^2] In fact, $\delta$ can be defined in a more intricate manner, such as incorporating the noise scale $\sigma$ as in (DBLP:conf/nips/Canonne0S20). However, this approach introduces a recursive dependency thus uncontrollable in practical applications, as $\sigma$ itself should be defined in terms of $\delta$. the term $\delta/\exp(\varepsilon)$ is a constant. Thus, it means that the distance between the distributions of $\mathcal{M}(x_{1})$ and $\mathcal{M}(x_{2})$ must be strictly bounded, i.e. similar to pure $\varepsilon$-LDP. However, this definition is often interpreted as: it is $\varepsilon$-LDP with a failure probability of $\delta$. Meanwhile, the Gaussian mechanism relies on this interpretation.
@@
Other relaxed privacy notions, including the Concentrated privacy from Dwork (DBLP:journals/corr/DuchiWJ16), Rényi privacy (DBLP:conf/csfw/Mironov17), and PAC privacy (DBLP:conf/crypto/XiaoD23), provide patches to $(\varepsilon,\delta)$-LDP. The Gaussian mechanism has natural results under the Concentrated privacy and Rényi privacy notions, while we provide a PAC privacy result in Appendix B.2. Among these privacy notions, $(\varepsilon,\delta)$-PAC LDP provides the same meaning as the common interpretation of $(\varepsilon,\delta)$-LDP: it is $\varepsilon$-LDP with a failure probability of $\delta$.
C.5. Other Discussions (Section 6)
C.5.1. Complexity of Finding the Robustness Radius
C.5.2. Per-instance Data Utility vs Aggregated Data Utility
Our utility quantification framework focuses on classifiers and per-instance data utility, providing utility guarantees at the level of individual data points. This stands in contrast to aggregated data utility analyses such as mean estimation, which evaluate the accuracy of aggregated statistics computed from the perturbed data of many users. Per-instance data utility is crucial for several reasons: (i) Classifiers and other personalized services like recommendation systems operate on individual data points. Their performance depends directly on per-instance utility and has no direct connection to aggregated utility. (ii) Per-instance data utility focuses on mechanism-level utility analysis, characterizing the privacy-utility tradeoff curves without relying on sample size or aggregation effects. These two reasons relate to our focus on per-instance data utility in this paper.
C.5.3. Fixed Classifier with Noisy Input vs Retrained Classifier with Noisy Data
These are two different scenarios. (i) In the former scenario, users locally apply an LDP mechanism to their data for privacy protection, agnostic to the service provider (i.e. the classifier), and are affected by the resulting utility of the existing classifier. (ii) In the latter scenario, the service provider applies LDP to users’ data and retrains a new classifier tailored to the perturbed data, which relates to “robust training”.
We focus on the former user-privacy-centered scenario; in the latter scenario, users expose their raw data to the service provider. Moreover, from users’ perspective, the classifier is fixed at the time of use (i.e. when they query the service). If a retrained classifier is deployed, then the classifier utility is for the retrained classifier (from users’ perspective).
C.5.4. Robustness of LDP Mechanisms
Intuitively, LDP mechanisms with higher concentration around raw data are more prone to reconstruction (with sufficient reports) and poisoning attacks, as malicious data are less averaged. In our concentration analysis, PM, SW, and $k$-$R$ (especially with small $k$) are more concentrated, thus expected to be more vulnerable. This aligns with prior findings: PM and $k$-$R$ show more vulnerability to poisoning attacks (DBLP:conf/uss/LiLSG023; DBLP:conf/ndss/LiLLS25).
C.5.5. Independent LDP Mechanisms vs Correlated LDP Mechanisms
We adopt independent perturbations for cleanness of presentation. In practice, DP libraries (e.g. Google DP (googledifferential-privacy_2026)) typically provide independent (L)DP primitives and do not explicitly support correlated perturbations. The essential reason is that the underlying data correlations are often unknown.
When the correlation coefficients are known or estimable, independent LDP can be extended via (i) post-processing of perturbed data to enforce a given correlation structure, or (ii) adopting correlated-perturbation mechanisms (e.g. (DBLP:journals/corr/abs-2507-17516; DBLP:conf/satml/VithanaCCJ25); although they are primarily tailored to mean estimation). Our utility quantification framework also applies to such correlated mechanisms by replacing the CDF of independent mechanisms with the conditional CDF of correlated mechanisms.
@@ C.5.5. Independent LDP Mechanisms vs Correlated LDP Mechanisms
Figure 9. Empirical and theoretical utility for two classifiers trained on the Stroke Prediction dataset.
(a) Logistic Regression (LR).
(b) Random Forest (RF).
Figure 10. Projected decision boundary of the two classifiers on the stroke dataset.
C.6. More Case Studies for the Stroke Prediction Dataset (Section 7.2.1)
C.6.1. Other Records and Mechanisms
@@
C.6.2. Closed Form CDF of the PM Mechanism
@@
C.6.3. Projected Decision Boundary
@@
(a) LR: Average-case utility.
(b) LR: Worst-case utility.
Figure 11. Average-case and worst-case utility for the Logistic Regression classifier trained on the Stroke Prediction dataset.
(a) RF: Average-case utility.
(b) RF: Worst-case utility.
Figure 12. Average-case and worst-case utility for the Random Forest classifier trained on the Stroke Prediction dataset.
C.6.4. Average-case and Worst-case Utility
@@
Conclusion on Robustness. From the results of average-case and worst-case utility analysis of the Logistic Regression and Random Forest classifiers, we can conclude that the Logistic Regression classifier is more robust under LDP-perturbed inputs than the Random Forest classifier.
C.7. More Case Studies for the Bank Customer Attrition Dataset (Section 7.2.2)
C.7.1. Detailed Quantification
@@
Figure 13. Average-case and worst-case utility for the Logistic Regression classifier trained on the Bank Customer Attrition dataset.
@@
Figure 14. Average-case and worst-case utility for the Random Forest classifier trained on the Bank Customer Attrition dataset.
C.7.2. Average-case and Worst-case Utility
C.8. More Details for the MNIST Classifier (Section 7.2.3)
C.8.1. Monte Carlo Estimation of the Utility
We emphasize the conceptual connection between LDP and robustness while adopting robustness notions that largely align with the standard local robustness used in prior work. Although these notions are clean and well established, they can yield conservative utility estimates in high-dimensional settings.
To obtain tighter utility quantification in high dimensions, one can consider more flexible robustness notions, e.g. robustness regions given by irregular sets rather than axis-aligned hyperrectangles (as induced by complex decision boundaries). Such regions can be estimated via Monte Carlo sampling in high-dimensional spaces. This estimation is performed once; afterward, for any given $\varepsilon$, the utility can be theoretically quantified by approximately integrating the $d$-dimensional LDP mechanism’s probability distribution over the estimated robustness region.
Formally, let $h:[0,1]^{d}\to\{1,\ldots,K\}$ be a $d$-dimensional classifier and let $x\in[0,1]^{d}$ be an input with predicted (and correct) label $h(x)$. Draw $\tilde{x}_{1},\ldots,\tilde{x}_{N}$ i.i.d. uniformly from $0,1]^{d}$, and for each sample check whether $h(\tilde{x}_{i})=h(x)$ holds. Assume that $m$ of these samples satisfy the check, and let $\mathcal{S}\subseteq[0,1]^{d}$ denote an estimated robustness region that contains the accepted samples. Under a $d$-dimensional (continuous) LDP mechanism, [^3] For discrete LDP mechanisms, the utility can be computed directly by summing the probabilities that the mechanism outputs samples in $\mathcal{S}$, eliminating the need for volume estimation. We omit this discussion for simplicity. the utility can then be quantified as
$\begin{split}\rho(\varepsilon,\mathcal{S})=&\int_{\mathcal{S}}pdf[\mathcal{M}_{\varepsilon}(x)=\tilde{x}]\mathrm{d}\tilde{x}\\ \approx&\mathrm{Vol}(\mathcal{S})\cdot\frac{1}{m}\sum_{i=1}^{m}pdf[\mathcal{M}_{\varepsilon}(x)=\tilde{x}_{i}],\end{split}$
where $\mathrm{Vol}(\mathcal{S})$ is the volume of the region $\mathcal{S}$, which can be estimated by $\mathrm{Vol}([0,1]^{d})\cdot m/N$. It is still a theoretical measure, since samples are fixed and $pdf[\mathcal{M}_{\varepsilon}(x)=\tilde{x}_{i}]$ is determined analytically by the LDP mechanism and can be evaluated for any $\varepsilon$.
This approach follows the connection between concentration analysis and robustness analysis established in this paper, with the concentration analysis now performed over the estimated robustness region $\mathcal{S}$ rather than the hyperrectangle $\theta_{\diamond}$. The error is now governed by the Monte Carlo estimation error (i.e. $\mathcal{O}(1/\sqrt{N})$) and volume estimation error, instead of Hoeffding bound in Theorem 6.
We experimented with this approach on the MNIST-$\times7$ dataset. While it is theoretically sound, in such a high-dimensional space the samples that pass the uniform-sampling check are sparse. As a result, for most accepted samples $\tilde{x}_{i}$, the density $pdf[\mathcal{M}_{\varepsilon}(x)=\tilde{x}_{i}]$ is close to zero, which drives the estimated utility to an unrealistically small value. Moreover, we cannot directly apply importance sampling using $pdf[\mathcal{M}_{\varepsilon}(x)]$ as the proposal distribution, since this proposal distribution depends on $\varepsilon$ and would essentially reduce the procedure to empirical utility estimation.
C.8.2. Average-case and Worst-case Utility
@@
Figure 15. Average-case and worst-case utility for the Neural Network classifier trained on the MNIST-$\times7$ dataset.
We use the first 50 images from the MNIST-$\times7$ dataset as representative samples for analyzing the average-case and worst-case utility under LDP-perturbed inputs. Figure 15 presents the results of this analysis. Both the average-case and worst-case utility show a similar trend to the point-wise utility quantification, with the PM mechanism consistently achieving the highest utility.